a) If the saving function is given by S° = 0.03Y^2 - 2Y + 150
Calculate the values of marginal propensity to save (MPS) and marginal propensity to
consume (MPC) when Y=40.

b) Find the second derivatives of these functions:
i) y=5x^2+350x-120 [3 marks]

ii) y= ln( x^2+4) [3 marks]

c) Find the relative maximum and minimum fory=3x^2+30x+605. [4 marks]
d) Find the absolute maximum and minimum for y=x^2-2x+3 between x=0 and x=3. [4 marks]
1
Expert's answer
2020-12-14T10:33:07-0500

a) Marginal propensity to save (MPS)

The marginal propensity to save (MPS) function is expressed as the derivative of the savings (S) function with respect to disposable income (Y).


"MPS=\dfrac{dS}{dY}=0.06Y-2"

When "Y=40"


"MPS=0.06(40)-2=0.4"

"1=MPC+MPS"

"MPC=1-MPS=1-(0.06Y-2)"


"=3-0.06Y"

When "Y=40"


"MPC=1-0.4=0.6"

When "Y=40, MPS=0.4, MPC=0.6"


b) Find the second derivatives of these functions:

i)

"y'=10x+350"

"y''=10"

ii)

"y'=\dfrac{2x}{x^2+4}"

"y''=2\cdot\dfrac{x^2+4-2x^2}{(x^2+4)^2}=-2\cdot\dfrac{x^2-4}{(x^2+4)^2}"


c)


"y'=6x+30"

Critical number(s)


"y'=0=>6x+30=0=>x=-5"

Critical number: "-5"


"y''=6>0"

"y(-5)=3(-5)^2+30(-5)+605=530"

By the Second DerivativeTest the function "y" has the relative minimum with the value of "530" at "x=-5."

The function "y" has no the relative maximum.


d)


"y'=2x-2"

Critical number(s)


"y'=0=>2x-2=0=>x=1"

Critical number: "1"

"y(0)=(0)^2-2(0)+3=3"

"y(3)=(3)^2-2(3)+3=6"

"y(1)=(1)^2-2(1)+3=2"


"2<3<6"

The function "y" has the absolute maximum with the value of "6" at "x=3" on "[0, 3]."

The function "y" has the absolute minimum with the value of "2" at "x=1" on "[0, 3]."